the area of a rectangular desk is given by the trinomial x2-7x-18.what are the possible dimensions of the desk?use factoring

Answer by ewatrrr(24785)

 
Hi,
the area of a rectangular desk is given by the trinomial x2-7x-18
A(rectangle) = l*w
x^2 -7x-18 = (x-9)(x+2) |(x-9)and (x+2) the possible dimensions
F First terms x^2
O Outside terms 2x
I Inside terms -9x
L Last terms -18